OpenStudy (anonymous):

Divide -28/41divided by (-7)

OpenStudy (anonymous):

-27/41/-7 = (-27) * (-7) / 41 = 189/41

Parth (parthkohli):

= -28/41 * 1/-7 = 4/41 * 1/1 = 4/41

Parth (parthkohli):

@SomeoneYouUsedToKnow Your second step is wrong, please check the answer again :)

OpenStudy (anonymous):

Oh thanks you are right @ParthKohli :)